OpenStudy (anonymous):

5m^2-11m-3=0 solve by using the quadratic formula

OpenStudy (anonymous):

i tried doing them myself and got them wrong so plzzz help me

OpenStudy (anonymous):

\[m^2-11m-3=0\] \[a=5,b=-11,c=-3\]

OpenStudy (anonymous):

formula is \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

we plug in the numbers and get \[\frac{-(-11)\pm\sqrt{(-11)^2-4\times 5\times (-3)}}{2\times 5}\]

OpenStudy (anonymous):

the roots are, m=(11+sqrt(181))/10 and m=(11-sqrt(181))/10

OpenStudy (anonymous):

now comes the computation \[\frac{11\pm\sqrt{121+60}}{10}\]

OpenStudy (anonymous):

\[\frac{11\pm\sqrt{181}}{10}\]

OpenStudy (anonymous):

and if i didn't make an arithmetic mistake that is it!